The density of air at 20°C and 1.00 atm is 1.205 g/L. If this air were compressed at the 

same temperature to equal the pressure at 40.0 m below sea level, what would be its 

density? Assume the barometric pressure is constant at 1.00 atm. The density of seawater 

is 1.025 g/cm3

.


1
Expert's answer
2021-06-03T05:44:32-0400

20 × 1 × 1.205 = 24.1

40/24.1 = 1 × 1.205 × X

1.66 = 1.205X

= 1.377g/L